# SkewRoute: Training-Free LLM Routing for Knowledge Graph Retrieval- Augmented Generation via Score Skewness of Retrieved Context

### v1

@@ A.3.2 Cumulative Frequency Distribution
Figure 11: Cumulative Frequency Distribution of Context Scores. (a)(b) Results on the WebQSP Dataset (c)(d) Results on the CWQ Dataset.
A.3.3 Correlation Analysis
To empirically validate the potential association between score skewness and query difficulty, we execute the following experimental analyses. Score skewness can be characterized using multiple metrics: area under scores, gini coefficient, entropy, and the number of contexts required to accumulate to a specified probability (e.g., 95%). High skewness typically has a smaller area, larger gini coefficient, lower entropy, and require fewer contexts for a high cumulative probability. Conversely, low skewness generally has a larger area, smaller gini, higher entropy, and require more contexts for the same threshold. Query difficulty is measured by the position of the correct answer within the descendingly sorted top-$K$ contexts. In KG-RAG, the answer position is a crucial indicator of retrieval quality. A lower answer position indicates the true answer is located at the head of the retrieved contexts, reflecting higher retrieval accuracy and facilitating LLMs’ generation. Conversely, late ranks may indicate insufficient discrimination of retriever or missing crucial context, increasing generation difficulty.
(a) Area
(b) Cumulative Threshold
(c) Entropy
(d) Gini Coefficient
@@ A.3.3 Correlation Analysis
Figure 12: Correlation Between Answer Position and Score Skewness. The subfigures show results for different skewness calculation metrics: (a) Area, (b) Cumulative Threshold, (c) Entropy, and (d) Gini Coefficient. Boxes span the 25th-75th percentiles with the median line. The red line shows the mean answer position of each group. ANOVA tests whether the mean answer positions differ significantly across the four groups.
The dataset is partitioned into four mutually exclusive groups based on each metric’s quartiles (Q1: 25th percentile, Q2: median, Q3: 75th percentile). The ANOVA indicates that mean answer positions differ significantly across the four groups. And we can clearly observe the trend in answer’s position in the retrieved contexts with score skewness. Specifically, as score skewness increases, the answer tends to appear at lower ranks (later positions) within the contexts. This indicates that simpler queries are associated with higher score skewness, whereas more difficult queries correlate with lower skewness, which confirms the strong correlation between query difficulty and the retrieved score skewness. Comparing area and gini coefficient (Figure 12(a) and Figure 12(d)), we find that the mean answer position varies more substantially with the gini coefficient than with area across groups, indicating that the gini coefficient offers greater sensitivity to score skewness. This suggests gini coefficient may more effectively to highlight difficulty of queries, which confirmed by our experimental section.
A.4 Prompts

### v2

@@ A.5 Cumulative Probability of Context Scores
Figure 7: Cumulative Probability of Context Scores. (a)(b) Results on the WebQSP Dataset (c)(d) Results on the CWQ Dataset.
A.6 Routing Across Cumulative Probability
(a) WebQSP Hit@1
(b) WebQSP F1-Score
(c) CWQ Hit@1
(d) CWQ F1-Score
A.7 Correlation Analysis
(a) WebQSP Dataset
(b) CWQ Dataset
Figure 9: Simple Query Number Across Skewness.
In the absence of supervision, there are no explicit labels to directly assess query difficulty. Nevertheless, in the RAG scenario, the retrieval quality directly impacts the answer generation of LLMs. The retrieval of clear and sufficient contexts inherently lowers the difficulty of answering the given query. A query is trivially simple if it can be answered with a single triple that appears right next to it in the prompt. Thus, to assess query difficulty, we use a straightforward metric: we check whether the top-ranked context in retrieved set contains the answer for given query. Simple query requires only a single knowledge context to answer. If the Top-$1$ contains real answer, we consider this query is simple. In this case, the scorer has accurately identified the context critical for reasoning with high confidence.
Next, we investigate the relationship between score skewness of retrieved contexts and query difficulty. To this end, we extract queries whose answers appear in the top-scoring context and compute the areas under their scores. Smaller areas correspond to higher skewness. We partition the area from their minimum to maximum into several contiguous intervals, each corresponding to a distinct level of skewness. Figure 9 illustrates the count of simple queries across different skewness levels. Overall, a consistent trend was observed in both datasets, which queries are predominantly concentrated at higher skewness levels. This allows us to select simple queries based on their skewness. The finding further reveals the strong correlation between query difficulty and score skewness, which can serve as effective indicator for routing LLMs to balance performance and cost.
(a) WebQSP Gini-based
(b) CWQ Gini-based
(c) WebQSP Entropy-based
(d) CWQ Entropy-based
(e) WebQSP Cumu-based
(f) CWQ Cumu-based
@@ A.7 Correlation Analysis
Figure 10: Correlation Between Query Difficulty and Score Skewness. The subfigures show results for different skewness calculation metrics on WebQSP and CWQ dataset: (a)(b) Gini Coefficient, (b)(d) Entropy, (e)(f) Cumulative Threshold. The red line shows the mean answer position of each group.
Furthermore, our gini-based, entropy-based and cumulative-based metrics for distribution skewness exhibit a trend consistent with the area-based method, as depicted in Figure 10. The rank of the answer-containing context can indicate query difficulty, where a lower position typically corresponds to a more complex, multi-hop reasoning scenario. We can observe a strong correlation where as the score distribution becomes more skewed, the rank of answer-containing context is higher, consequently, lower query difficulty. This reinforces the core insight: score skewness is a reliable indicator of query difficulty, enabling a dynamic routing strategy to balance inference performance and cost across different LLMs.
Dataset Method 0% 20% 40% 60% 80% 100% Avg. Eff.
WebQSP Random Routing 77.52 78.18 78.85 79.51 80.18 80.84 -
RouteLLM [ICLR, 25] 77.52 78.56 (+0.38) 79.18 (+0.33) 79.55 (+0.04) 79.98 (-0.20) 80.84 +0.14
GraphRouter [ICLR, 25] 77.52 78.26 (+0.08) 79.30 (+0.45) 80.22 (+0.71) 80.71 (+0.53) 80.84 +0.44
Ours (Area-based) 77.52 78.26 (+0.08) 80.10 (+1.25) 81.08 (+1.57) 81.14 (+0.96) 80.84 +0.97
Ours (Gini-based) 77.52 79.48 (+1.30) 79.98 (+1.13) 81.20 (+1.69) 80.96 (+0.78) 80.84 +1.23
Ours (Entropy-based) 77.52 79.18 (+1.00) 80.34 (+1.49) 81.08 (+1.57) 80.77 (+0.59) 80.84 +1.16
Ours (Cumulative-based) 77.52 79.24 (+1.06) 80.04 (+1.19) 80.71 (+1.20) 80.71 (+0.53) 80.84 +1.20
CWQ Random Routing 45.68 47.59 49.51 51.42 53.34 55.25 -
RouteLLM [ICLR, 25] 45.68 47.81 (+0.22) 50.41 (+0.90) 52.00 (+0.58) 53.61 (+0.27) 55.25 +0.49
GraphRouter [ICLR, 25] 45.68 47.83 (+0.24) 50.10 (+0.59) 52.02 (+0.60) 53.78 (+0.44) 55.25 +0.47
Ours (Area-based) 45.68 48.06 (+0.47) 49.59 (+0.08) 52.02 (+0.60) 53.95 (+0.61) 55.25 +0.44
Ours (Gini-based) 45.68 48.94 (+1.35) 50.92 (+1.41) 52.53 (+1.11) 54.15 (+0.81) 55.25 +1.17
Ours (Entropy-based) 45.68 48.74 (+1.15) 50.72 (+1.21) 52.70 (+1.28) 54.01 (+0.67) 55.25 +1.08
Ours (Cumulative-based) 45.68 49.02 (+1.43) 50.89 (+1.38) 52.82 (+1.40) 53.89 (+0.55) 55.25 +1.19
Table 11: Routing Between Qwen2.5 Models: 7B as the Small LLM and 72B as the Large LLM with Area-based Routing Method. The best scores are highlighted with bold, the second-best scores are highlighted with underline and the third-best scores are indicated in italics.
Dataset Method 0% 20% 40% 60% 80% 100% Avg. Eff.
WebQSP Random Routing 78.56 79.68 80.80 81.91 83.03 84.15 -
RouteLLM [ICLR, 25] 78.56 79.73 (+0.05) 80.47 (-0.33) 81.82 (-0.09) 82.86 (-0.17) 84.15 -0.14
GraphRouter [ICLR, 25] 78.56 79.67 (-0.01) 80.65 (-0.15) 81.39 (-0.52) 83.35 (+0.32) 84.15 -0.09
Ours (Area-based) 77.52 80.22 (+0.54) 81.76 (+0.96) 82.43 (+0.52) 83.91 (+0.88) 80.84 +0.73
Ours (Gini-based) 78.56 81.33 (+1.65) 81.57 (+0.77) 82.62 (+0.71) 83.35 (+0.32) 84.15 +0.86
Ours (Entropy-based) 78.56 81.08 (+1.40) 82.00 (+1.20) 82.49 (+0.58) 83.66 (+0.63) 84.15 +0.95
Ours (Cumulative-based) 78.56 81.57 (+1.89) 81.27 (+0.47) 82.62 (+0.71) 83.60 (+0.57) 84.15 +0.91
CWQ Random Routing 49.90 51.51 53.12 54.72 56.33 57.94 -
RouteLLM [ICLR, 25] 49.90 51.26 (-0.25) 53.61 (+0.49) 55.03 (+0.31) 56.75 (+0.42) 57.94 +0.24
GraphRouter [ICLR, 25] 49.90 50.84 (-0.67) 52.79 (-0.33) 54.80 (+0.08) 56.16 (-0.17) 57.94 -0.27
Ours (Area-based) 49.90 52.25 (+0.74) 53.87 (+0.75) 55.45 (+0.73) 56.73 (+0.40) 57.94 +0.66
Ours (Gini-based) 49.90 52.65 (+1.14) 55.00 (+1.88) 56.16 (+1.44) 57.04 (+0.71) 57.94 +1.29
Ours (Entropy-based) 49.90 52.51 (+1.00) 54.89 (+1.77) 56.07 (+1.35) 56.78 (+0.45) 57.94 +1.14
Ours (Cumulative-based) 49.90 52.68 (+1.17) 54.77 (+1.65) 56.41 (+1.69) 56.61 (+0.28) 57.94 +1.20
Table 12: Routing Between Llama3.1 Models: 8B as the Small LLM and 70B as the Large LLM with Area-based Routing Method. The best scores are highlighted with bold and the second-best scores are highlighted with underline and the third-best scores are indicated in italics.
A.8 Results of Area-based Routing Method
To evaluate the area-based routing method, we conduct experiments on the WebQSP and CWQ datasets using Qwen2.5-7B-Instruct and Qwen2.5-72B-Instruct. The area under the context score curve naturally reflects skewness. High skewness corresponds to a rapid decline, yielding a smaller area enclosed with the coordinate axes, whereas low skewness indicates gradual declines with more high-scoring contexts, thus forming a larger area. Despite being simple and intuitive, this area-based method proves effective, as demonstrated by Table 11. Overall, the area-based method consistently outperforms all baselines on the WebQSP dataset. Specifically, although slightly less competitive than other three methods, it still surpasses RouteLLM nearly 7x and GraphRouter over 2x. On the CWQ dataset, the area-based method performs on par with RouteLLM and GraphRouter, while requiring no training. These findings underscore the potential of area-based method as an effective and training-free solution for query routing in knowledge-intensive tasks.
Table 12 presents the routing results between Llama3.1-8B-Instruct and Llama3.1-70B-Instruct. The area-based method demonstrates strong plug-and-play generalization, requiring no training while still delivering consistent gains over baseline routers. On the WebQSP dataset, the area-based method achieves an Average Effectiveness of 0.73, even when RouteLLM and GraphRouter fail. On the CWQ dataset, the area-based method remains effective, outperforming the strong baseline RouteLLM by nearly 3x. These results highlight the area-based method as a training-free, plug-and-play and readily deployable routing strategy with strong generalization across both datasets and model families.
A.9 Prompts